12,780 44,290 48,150 WIP beginning (30% for conversion) Units started Units completed and transferred out WIP ending (40% for conversion) Beginning WIP direct materials Beginning WIP conversion costs Cost of materials added Cost of conversion added 23,940 8,820 $ 192,926 $ 172,193 What are the total physical units to account for? How many units are in ending WIP?[ What are the equivalent units of production for materials? What are the equivalent units of production in ending WIP for conversion? What are the total equivalent units of production for conversion? What are the total costs to account for? What is the cost per equivalent unit for materials? What is the cost per equivalent unit for conversion?[ What is the total cost of the units completed and transferred out? What is the total cost of the units in ending WIP?
Answer #1

Answer 1:

Workings:

WIP Beginning Units started Total Physical units to account 57,070 12,780 44,290

Answer 2:

Units in Ending WIP 8,920

Workings:

WIP Beginning Units started Total Physical units to account for Less, Units completed and transferred out 48,150 Units in Ending WIP 12,780 44,290 57,070 8,920

Answer 3:

Equivalent units of production for material 57,070

Workings:

Weighted Average cost method: Material Conversion Equivalent Units: Units completed and transferred out Ending WIP 48,150 48,150 Material 100% Conversion 40% 8,920 3,568 51,718 Equivalent units of production 57,070

Answer 4:

Please workings in answer 3 above

Answer 5:

Please workings in answer 3 above

Answer 6:

Total cost to account for$397,879

Workings:

Material Conversion Total Cost: WIP Beginning Added during the period Total ,940 $8,820 $32.760 $192,926 $172,193 $365,119 $216,866 $181,013 $397,879

Answer 7:

Workings:

Cost per equivalent unit for material = Total Cost for material / EUP for material = $216,866 / 57,070 = $3.80

Answer 8:

Workings:

Cost per equivalent unit for material = Total Cost for conversion / EUP for conversion = $181,013 / 51,718 = $3.50

Answer 9:

Workings:

Total cost of units completed and transferred out = ($3.80 + $3.50) * 48,150 = $351,495

Answer 10:

Workings:

Total cost of units in ending WIP = $3.80 * 8920 + $3.50 * 3568 = $46,384